OpenStudy (anonymous):

In the solution of the equation 5 - 3x = 2x + 9, 3x is added to the equation first. Which of the following should be done next? add -9 add -5 add -2x divide by 5

OpenStudy (skullpatrol):

Any ideas?

mathslover (mathslover):

I think it should be i) add -9 as we have : 5 - 3x = 2x + 9 5 - 3x + 3x = 2x + 3x +9 5 = 5x + 9 now to proceed we must remove 9 from RHS so that we can get the answer easily. So i) add -9 is the best choice

mathslover (mathslover):

Did it help you @brookellerenee ?